OpenStudy (anonymous):

I need help with multiplying fractions i will post the fraction in the reply section will medal and fan

OpenStudy (anonymous):

\[-\frac{ 5}{ 6}\div \frac{ 9 }{ 10}\]

OpenStudy (pawanyadav):

What you think

OpenStudy (anonymous):

i know you keep the first fraction the same, ,change division to multiplication, and replace the 2nd fraction with its reciprocal

OpenStudy (pawanyadav):

Than what you need help in?

OpenStudy (anonymous):

so then your equation will be\[-\frac{ 5}{ 6}\times \frac{ 9 }{ 10}\]

OpenStudy (anonymous):

i dont know if i multiply the 2 numbers in the first fraction as negatives

OpenStudy (pawanyadav):

reciprocal it to10/9

OpenStudy (anonymous):

oh yea forgot about that XD

OpenStudy (anonymous):

so when i do that do i do -5*9 and -6*10

OpenStudy (pawanyadav):

You will get. -5/6 × 10/9 So simply multiply themthem -5*10/(6*9)=?

OpenStudy (pawanyadav):

Do you get that?

OpenStudy (anonymous):

okay so i should get -50/54 simplified to -25/27

OpenStudy (pawanyadav):

Yes

OpenStudy (anonymous):

okay thanks!

OpenStudy (pawanyadav):

Welcome